OpenStudy (anonymous):

Given f(x)=−4x^2−7x, determine the points where (a) the curve intercepts the x− axis; (b) the tangent line to the curve is horizontal.

OpenStudy (anonymous):

(a) f(x) = 0 (b) df(x)/dx = 0

OpenStudy (anonymous):

i'm sorry can u explain? it?

OpenStudy (anonymous):

what can you tell me about the x axis? what is the value of y on the x axis?

OpenStudy (anonymous):



OpenStudy (anonymous):

zero is the value.

OpenStudy (anonymous):

right. so at the point that the curve intercepts the x axis, what is the value of f(x) ?

OpenStudy (anonymous):

0? is it?

OpenStudy (anonymous):

but how do we find the value for x?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

if f(x) = 0, that means-4x^2 - 7x = 0 since f(x) = -4x^2 -7x

OpenStudy (anonymous):

so x=0 or x=7/-4?

OpenStudy (anonymous):

you have to do the calculations yourself. I am not going to do your calculations for you. plug your equation into the roots for ax^2 + bx +c = 0

OpenStudy (anonymous):

why? cant we just factor (-4x^2-7x) and get x(-4x-7) then find the value for x?

OpenStudy (anonymous):

sure, that works too.

OpenStudy (anonymous):

so, what the points where the curve intercepts the x axis?

OpenStudy (anonymous):

is it (7/-4,0)

OpenStudy (anonymous):

but i dont think thats right.

OpenStudy (anonymous):

why not? also, you are missing one more point

OpenStudy (anonymous):

(0,0) is the other one

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok i get it! so for me we just plug in 0=-8x-7?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you mean for the second part? yes,

OpenStudy (anonymous):

yes i meant to say for "b)"

OpenStudy (anonymous):

Thank you! so much!

OpenStudy (anonymous):

happy to help